RC meeting and Run Plan, Thursday Oct. 4

From HallCWiki
Jump to: navigation, search
  • Status
    • Kaon LT status page
    • Beam energy measurement done [1]
    • Harp scans a the target after beam energy measurement looked good [2]
    • Luminosity scans deferred to a later time so we can take the very bare minimum statistics for the current kinematic point

  • Issues
    • Need higher beam delivery eff. since we are sooo behind

  • Plans - Physics by Friday 7 am
    • Finish the "right" kinematics then go back to "center" and finish taking 25 K. If time left, take more statistics for "right", "center", "left"

  • Plans - Access Friday
    • RadCon will come in for survey Friday at 7am
    • SHMS aerogel tray change: ~6-8 hours; For this we have to:
        • move HMS to 38 deg (to accommodate the roof of SHMS on the floor)
        • turn off SHMS detectors HVs except for S1Y5B
        • place the plastic cover on top of the detector stack
        • use the HGC window protection thingy
    • Jack, Steve will go in to check on that HMS Dipole LN2 JT valve
    • Howard will go in to check on the over-active group of wires in HMS DC1X2
    • 3HC07A harp checkout??
    • some other folks to look at the dump diffuser?
    • students will go in to work on the SoLID Cher setup